# VURF: A General-purpose Reasoning and Self-refinement Framework for Video Understanding

### v2

@@ Abstract
Recent studies have demonstrated the effectiveness of Large Language Models (LLMs) as reasoning modules that can deconstruct complex tasks into more manageable sub-tasks, particularly when applied to visual reasoning tasks for images. In contrast, this paper introduces a Video Understanding and Reasoning Framework (VURF) based on the reasoning power of LLMs. Ours is a novel approach to extend the utility of LLMs in the context of video tasks, leveraging their capacity to generalize from minimal input and output demonstrations within a contextual framework. By presenting LLMs with pairs of instructions and their corresponding high-level programs, we harness their contextual learning capabilities to generate executable visual programs for video understanding. To enhance program’s accuracy and robustness, we implement two important strategies. Firstly, we employ a feedback-generation approach, powered by GPT-3.5, to rectify errors in programs utilizing unsupported functions. Secondly, taking motivation from recent works on self refinement of LLM outputs, we introduce an iterative procedure for improving the quality of the in-context examples by aligning the initial outputs to the outputs that would have been generated had the LLM not been bound by the structure of the in-context examples. Our results on several video-specific tasks, including visual QA, video anticipation, pose estimation and multi-video QA illustrate the efficacy of these enhancements in improving the performance of visual programming approaches for video tasks. https://github.com/ahmad-573/VURF
Keywords:
Video Understanding Reasoning Vision Language
@@ 1 Introduction
Figure 1: An Overview of the VURF pipeline: Figure demonstrates how a complex query regarding video editing is broken down in our Video Programming framework to sequentially arrive at the final edited result. Best viewed in zoom.
@@
In recent years, the vision community has developed highly efficient specialized models for various video understanding tasks, including Video Question Answering [4], Action Anticipation [23] and Pose Estimation [58, 31, 70, 21]. Despite such advancements, the video models usually offer isolated visual comprehension capabilities in the form of narrow task-specific models. Such specialized models limited to individual tasks struggle to offer a comprehensive, adaptable, and scalable understanding of videos for complex reasoning. Although the current off-the-shelf models can perform specific well-defined tasks, they require specialized and comprehensive datasets to effectively train dedicated models for each task, which is not feasible for general-purpose complex reasoning problems. Furthermore, individual off-the-shelf vision models for each video-understanding task necessitate a distinct framework with unique model configurations. This problem underscores the requirement for a uniform reasoning framework that offers plug-and-play architecture, capable of leveraging any pre-trained computer vision model, enabling seamless execution of a given task.
@@
To address these challenges, we adopt a strategy of decomposing broad video-understanding tasks into more manageable sub-tasks, each of which can be solved by executing task-specific models, and subsequently consolidating the results. Our framework is motivated by the observation that complex tasks can be effectively solved by executing an intermediate sequence of sub-tasks, collectively working in sequence to solve the more challenging problem. This process of task decomposition necessitates a reasoning module capable of discriminating the necessary steps for task execution. Large Language Models (LLMs) emerge as promising candidates for this role. Recent works on Visual Programming [27] has demonstrated the effectiveness of LLMs in converting complex tasks into smaller, more tractable components that can be addressed by specialized computer vision models [64, 69, 47]. In our study, we demonstrate the utility of such a reasoning module in tackling specific challenges within the domain of video understanding. We also show that building such an approach on top of the existing off-the-shelf video models can significantly enhance task performance.
@@
While LLMs demonstrate competence in serving as reasoning modules, they are not immune to errors and limitations. One notable deficiency is their vulnerability to hallucinations induced by contextual information, without the means to self-correct based on task-agnostic knowledge [54]. To address this concern, we draw inspiration from recent research demonstrating the efficacy of self-refinement processes in enabling LLM’s to enhance their outputs [44, 19], akin to the way humans engage in self-correction. Specifically, we propose a feedback-generation mechanism that evaluates the LLM’s output and employs this feedback to prompt the LLM to refine its output. Our findings substantiate the effectiveness of this approach in elevating the performance of video-based reasoning approaches. Our main contributions are as follows:
@@
  • The first generic visual reasoning framework for video understanding consolidating multiple task-specific, domain-specialized video models to answer any video-related user queries.
@@
  • Using in-context learning, we align LLM behavior for decomposing a given complex task into multiple subtasks easily solvable using existing task-specific video models.
@@
  • Our proposed self-refinement strategy helps avoid errors in the programs (outlining subtask decomposition) generated by the LLM and boosts the performance by iteratively refining the generated program. The proposed framework is shown to boost performance for tasks such as visual question answering for videos in complex reasoning scenarios.
@@ 2 Related Work
Video Understanding: Video understanding focuses on teaching machines to understand and analyze visual content. One crucial task is to recognize and localize different actions in the video [46]. Numerous algorithms have been developed to cater the video understanding tasks such as Video Swin Transformer [41], VideoMAE [57], C2D [61], MViT-V2 [38], STGCN++ [71], and ViViT [5] that achieve high accuracy on SOTA datasets [51, 22, 26, 15, 28, 23, 48].
@@
Video understanding tasks facilitate the efficient handling of diverse information modalities [36] and various tasks have been introduced to test the capabilities of the methods for video understanding such as retrieving temporal and spatial information [76] and answering natural language questions from the video i.e., Video Question Answering (VQA) [68, 34]. Other methods such as SeViLA [72] and iVQA [40], adapt Localizer and Answerer for both QA and temporal key-frame localization which are then extended to zero-shot VQA [52, 67, 39]. Another challenging task in video understanding is to localize the starting and ending time of the video segment that corresponds to the input query i.e., video grounding [11, 67, 75]. It can solve various video understanding tasks such as Temporal Action Recognition [12], spatio-temporal video grounding [77], and Action Recognition [9]. Our work on video programming provides a pipeline for various video understanding and reasoning tasks leveraging off-the-shelf SOTA models and the reasoning capabilities of LLMs for each sub-task.
@@
Visual Programming: Visual Programming leverages the capabilities of Language Models (LLMs) as reasoning modules to break down complex vision-understanding tasks into simpler intermediate sub-tasks executed sequentially. Utilizing in-context examples along with prompts can effectively decompose these tasks, enhancing the responses generated by language models. In recent advancements within visual programming, such as VisProg [27], strategies have been proposed to enhance performance in visual tasks. These strategies include increasing the number of in-context examples, substituting high-error modules with recent off-the-shelf models, and fine-tuning instructions to achieve improved performance. Our work, VURF, stands out as the first generic reasoning framework for the video modality. It not only utilizes the self-critique strategy of LLMs but also employs periodic refinement of visual programs. This process further ensures the replacement of invalid functions with a valid module list, addressing misalignment errors effectively.
The concept of Visual Programming can be extended to any family of vision-language tasks such as BERT, and it’s inspired vision-language work [20, 37, 80]. Recently, zero-shot models such as Flamingo have emerged that do not require fine-tuning on new tasks [2]. These body of works solve different vision-understanding tasks such as Visual Question Answering, VLM for few-shot learning [2], Grounding, Segmentation, and Tracking. Moreover, the vision-language model incorporates multiple types of information and is trained using a contrasting approach which has made it possible to adapt to new tasks without the necessity of fine-tuning, allowing for zero-shot adaptation [29]. However, zero-shot models struggle to generalize and outperform SOTA in the unseen data as they lack adaptation on specific tasks [33, 60]. In contrast, VURF, leverages SOTA off-the-shelf models for the downstream tasks by decomposing complex tasks into visual programs that uses SOTA models.
@@
Self Refinement: LLMs demonstrate a special ability to enhance their outputs just like how humans re-evaluate, refine, and reiterate the text that they have initially written. The same LLM when used to identify the potential issues with the output and through the generator, refiner, and feedback provider, it even has the potential to improve responses generated by SOTA LLMs like GPT4 [43]. Other methods include using external tools like search engines to rectify the output [43], detecting hallucinated outputs [25, 18, 79, 24], using natural language feedback [49] to improve the initial generated response.
@@
Leveraging the capabilities of self-debugging, language models have been able to debug their predicted program via few-shot demonstrations [13], use of relevant in-context examples to generate efficient response [59, 56], and have contributed in various domains such as code generation and its applications [73], summarization [7], and program synthesis [32, 30].
@@
While these approaches improve the LLM’s response via recursive-feedback methods, these methods require continuous refinement of any output generated. Our self-refinement approach focuses on the critique and refinement of in-context examples and we show that just a pre-refined set of in-context examples can boost the performance of a visual programming approach (Table 1).
@@
Figure 2: Video Understanding and Reasoning Framework (VURF) pipeline. Top: figure shows the main approach of visual programming for videos with the added self-correction module. Bottom: figure shows the self-refinement module. Our proposed VURF framework allows improved reasoning for video understanding building on the existing individual task-specific models
@@ 3 Methodology
In contrast to conventional task-specific models that exhibit limitations in addressing complex reasoning challenges, the Video Understanding and Reasoning Framework (VURF) seeks to utilize the reasoning power of LLMs to deconstruct complex video-related queries into a series of sub-tasks (video programs). By executing these sub-tasks sequentially, we can culminate to arrive at the final response. Moreover, VURF allows seamless integration of new visual models in the plug-and-play manner and also employs the self-critique mechanisms to mitigate LLM’s hallucinations and judgement errors. An overview of our approach is shown in Fig. 2, and we explain the approach in detail below.
@@ 3.1 Video Reasoning LLMs
Role of Large Language Models. Large language models (LLMs), exemplified by GPT-3 [6] and GPT-4 [1], have demonstrated an impressive in-context learning capability (ICL) that does not necessitate fine-tuning. ICL aims to extend the understanding of LLMs to novel scenarios using a restricted set of input and output demonstrations within the relevant context. In this work, we leverage GPT-3.5 to generate visual programs that solves video reasoning tasks involving natural language instructions. The LLMs efficiently performs video reasoning tasks by avoiding direct video processing. Instead, it formulates logical sequence flows that decompose complex tasks into simpler sub-tasks. This approach enhances efficiency and allows the model to navigate the intricacies of video-related challenges.
Prompting. We prompt GPT-3.5 with in-context examples which consists of pairs of instruction and the associated programs that the LLM is expected to generate. The programs follow a generic structure where each line of the program includes the name of a module, the module’s input argument names and their values, and an output variable name. As output variables in a specific step are used later for another step they follow a general structure which the LLM learns:
@@ 3.2 Self-Refinement
One prominent limitation of a naive visual programming technique like [27] is the proneness of generating inaccurate information influenced by contextual cues, as well as lacking an inherent capacity for autonomous self-correction through task-agnostic knowledge. There are two major issues that arise with the LLM-generated program. Firstly, due to LLM hallucinations, the program might be using some function that is not supported by our interpreter. Secondly, the program may not break down the instruction into sub-tasks in an optimal fashion. We resolve these limitations by the below steps.
@@
Auto-Refinement of In-Context Examples: The effectiveness of our approach hinges significantly on the quality of in-context examples. Therefore, it is imperative to refine these examples to enhance the module’s performance [42, 43]. To accomplish this, we implement a self-refinement procedure. Initially, given an instruction, $I$, and the initial program generated $P$, we input $I$ to the LLM, prompting it to generate an improved program, $P^{\prime}$, without the inclusion of in-context examples. Subsequently, both $P$ and $P^{\prime}$ are input into the LLM, enabling the generation of a refined program that aligns more closely with the LLM’s reasoning, all while excluding the influence of in-context examples. Replicating this process for $n$ instructions yields a set of $n$ new in-context examples, enhancing the module’s ability to perform tasks effectively. Note that as shown in Fig. 2, the auto self-refinement module can be applied to a single user query to iteratively improve the generated program as well. However, this is inefficient and costly and thus we show that even pre-refining the in-context program-instruction pairs can improve the performance of the VURF (Table 1). A concrete illustration of this process on a single program and instruction pair is provided in Fig. 3.
Figure 3: Auto Self-Refinement demonstrated through an example. Two programs are generated: one is provided with contextual examples, and the other one is not - although certain information is appended to ensure the program’s structural integrity. Subsequently, both programs are input into the Language Model (LLM), with the task of generating a new program that closely aligns with the ideal program but does not use any invalid functions.
Manual Refinement of In-Context Examples: While the auto self-refinement approach enhances contextual relevance, it falls short in augmenting the functionality of the in-context examples. To overcome this limitation, we propose a methodology that involves leveraging the capabilities of the Large Language Model (LLM) to introduce new functions or modules (Fig. 4), thereby elevating the quality of our in-context examples. Our approach entails generating programs from a set of instructions and tasking the LLM with enhancing the program, without imposing constraints on the functions it may introduce. Through careful observation of the resulting program, we can identify and incorporate valuable additions to our list of functions.
@@ 4 Tasks
Our video understanding and reasoning framework is an effort to provide a comprehensive approach that can be easily adapted to various visual tasks by incorporating an interpreter component for an existing pre-trained SOTA vision model. VURF solves four diverse and comprehensive tasks, including Video Question Answering (VQA), Video Anticipation, Pose Estimation, and Multi-Video VQA. The program works by using LLM as a reasoning module to generate a visual program that consists of subsequent steps that are executed in isolation and the output of each predecessor as fed to the input to the successor.

### v3

@@ Abstract
Recent studies have demonstrated the effectiveness of Large Language Models (LLMs) as reasoning modules that can deconstruct complex tasks into more manageable sub-tasks, particularly when applied to visual reasoning tasks for images. In contrast, this paper introduces a Video Understanding and Reasoning Framework (VURF) based on the reasoning power of LLMs. Ours is a novel approach to extend the utility of LLMs in the context of video tasks, leveraging their capacity to generalize from minimal input and output demonstrations within a contextual framework. We harness their contextual learning capabilities by presenting LLMs with pairs of instructions and their corresponding high-level programs to generate executable visual programs for video understanding. To enhance the program’s accuracy and robustness, we implement two important strategies. Firstly, we employ a feedback-generation approach, powered by GPT-3.5, to rectify errors in programs utilizing unsupported functions. Secondly, taking motivation from recent works on self-refinement of LLM outputs, we introduce an iterative procedure for improving the quality of the in-context examples by aligning the initial outputs to the outputs that would have been generated had the LLM not been bound by the structure of the in-context examples. Our results on several video-specific tasks, including visual QA, video anticipation, pose estimation, and multi-video QA, illustrate these enhancements’ efficacy in improving the performance of visual programming approaches for video tasks.
@@ 1 Introduction
In recent years, the vision community has developed highly efficient specialized models for various video understanding tasks, including Video Question Answering Antol et al. (2015), Action Anticipation Girdhar and Grauman (2021) and Pose Estimation Toshev and Szegedy (2014); Koprinska and Carrato (2001); Yilmaz et al. (2006); Gammulle et al. (2019). Despite such advancements, video models usually offer isolated visual comprehension capabilities in the form of narrow task-specific models. Such specialized models limited to individual tasks struggle to offer a comprehensive, adaptable, and scalable understanding of videos for complex reasoning. Although the current off-the-shelf models can perform specific well-defined tasks, they require specialized and comprehensive datasets to effectively train dedicated models for each task, which is not feasible for general-purpose complex reasoning problems. Furthermore, individual off-the-shelf vision models for each video-understanding task necessitate a distinct framework with unique model configurations. This problem underscores the requirement for a uniform reasoning framework that offers plug-and-play architecture, capable of leveraging any pre-trained computer vision model, enabling seamless execution of a given task.
@@
To address these challenges, we adopt a strategy of decomposing broad video-understanding tasks into more manageable sub-tasks, each of which can be solved by executing task-specific models, and subsequently consolidating the results. Our framework is motivated by the observation that complex tasks can be effectively solved by executing an intermediate sequence of sub-tasks, collectively working in sequence to solve the more challenging problem. This process of task decomposition necessitates a reasoning module capable of discriminating the necessary steps for task execution. Large Language Models (LLMs) emerge as promising candidates for this role. Recent works on Visual Programming Gupta and Kembhavi (2023) have demonstrated the effectiveness of LLMs in breaking down complex tasks into smaller, more manageable components that can be tackled by specialized computer vision models Wu et al. (2023); Yang et al. (2023); Ruan et al. (2023). In our study, we demonstrate the utility of such a reasoning module in tackling specific challenges within the domain of video understanding. We also show that building such an approach on top of the existing off-the-shelf video models can significantly enhance task performance.
@@
While LLMs demonstrate competence in serving as reasoning modules, they are not immune to errors and limitations. One notable deficiency is their vulnerability to hallucinations induced by contextual information, without the means to self-correct based on task-agnostic knowledge Sun et al. (2023). To address this concern, we draw inspiration from recent research demonstrating the efficacy of self-refinement processes in enabling LLMs to enhance their outputs Madaan et al. (2024); Feng et al. (2024), akin to the way humans engage in self-correction. Specifically, we propose a feedback-generation mechanism that evaluates the LLM’s output and employs this feedback to prompt the LLM to refine its output. Our findings substantiate the effectiveness of this approach in elevating the performance of video-based reasoning approaches. Our main contributions are as follows:
@@
  • The first generic visual reasoning framework for video understanding consolidates multiple task-specific, domain-specialized video models to answer any video-related user queries.
@@
  • Using in-context learning, we align LLM behavior for decomposing a given complex task into multiple sub-tasks easily solvable using existing task-specific video models.
@@
  • Our proposed self-refinement strategy helps avoid errors in the programs (outlining subtask decomposition) generated by the LLM and boosts the performance by iterative refining the generated program. The proposed framework is shown to boost performance for tasks such as visual question answering for videos in complex reasoning scenarios.
Figure 1: An overview of the VURF pipeline: Figure demonstrates how a complex query regarding video editing is broken down in VURF to arrive at the final edited result. Best viewed in zoom.
@@ 2 Related Work
Video Understanding: Video understanding focuses on teaching machines to understand and analyze visual content. One crucial task is to recognize and localize different actions in the video Reddy et al. (2023). Numerous algorithms have been developed to cater the video understanding tasks such as Video Swin Transformer Liu et al. (2022), VideoMAE Tong et al. (2022), C2D Wang et al. (2018), MViT-V2 Li et al. (2022), STGCN++ Yu et al. (2017), and ViViT Arnab et al. (2021) that achieve high accuracy on SOTA datasets Sigurdsson et al. (2016); Gao et al. (2017); Gu et al. (2018); Deliege et al. (2021); Huang et al. (2020); Girdhar and Grauman (2021); Sadhu et al. (2021).
@@
Video understanding tasks facilitate the efficient handling of diverse information modalities Li et al. (2020) and various tasks have been introduced to test the capabilities of the methods for video understanding such as retrieving temporal and spatial information Zhang et al. (2023) and answering natural language questions from the video i.e., Video Question Answering (VQA) Yang et al. (2003); Lei et al. (2018). Other methods such as SeViLA Yu et al. (2024) and iVQA Lin et al. (2023b), adapt Localizer and Answerer for both QA and temporal key-frame localization which are then extended to zero-shot VQA Song et al. (2023); Yang et al. (2022); Lin et al. (2023a). Another challenging task in video understanding is to localize the starting and ending time of the video segment that corresponds to the input query i.e., video grounding Chen et al. (2018); Yang et al. (2022); Zeng et al. (2020). It can solve various video understanding tasks such as Temporal Action Recognition Chen et al. (2019), spatio-temporal video grounding Zhang et al. (2020), and Action Recognition Carreira and Zisserman (2017). Our work on video programming provides a pipeline for various video understanding and reasoning tasks leveraging off-the-shelf SOTA models and the reasoning capabilities of LLMs for each sub-task.
@@
Visual Programming: Visual Programming leverages Language Models (LLMs) to break down complex vision-understanding tasks into simpler sub-tasks executed sequentially, improving responses with in-context examples and prompts. Recent advancements, such as VisProg Gupta and Kembhavi (2023), enhance visual task performance by increasing in-context examples, replacing high-error modules with off-the-shelf models, and refining instructions. Our work, VURF, is the first generic reasoning framework for video, utilizing LLMs’ self-critique and refining visual programs to address errors. While zero-shot models like Flamingo Alayrac et al. (2022) can adapt to new tasks without fine-tuning, they struggle with generalization, unlike VURF, which leverages SOTA models for downstream tasks.
@@
Self Refinement: LLMs demonstrate a special ability to enhance their outputs just like how humans re-evaluate, refine, and reiterate the text that they have initially written. The same LLM when used to identify the potential issues with the output and through the generator, refiner, and feedback provider, it even has the potential to improve responses generated by SOTA LLMs like GPT4 Madaan et al. (2023). Other methods include using external tools like search engines to rectify the output Madaan et al. (2023), detecting hallucinated outputs Gou et al. (2023); Evans et al. (2021); Zhou et al. (2020); Golovneva et al. (2022), using natural language feedback Saunders et al. (2022) to improve the initial generated response.
@@
Leveraging the capabilities of self-debugging, language models have been able to debug their predicted program via few-shot demonstrations Chen et al. (2023), use of relevant in-context examples to generate efficient response Wang et al. (2023); Thawakar et al. (2024), and have contributed in various domains such as code generation and its applications Yu et al. (2019), summarization Campos and Shern (2022), and program synthesis Le et al. (2022); Kim et al. (2023).
@@
While these approaches improve the LLM’s response via recursive feedback methods, these methods require continuous refinement of any output generated. Our self-refinement approach focuses on the critique and refinement of in-context examples and we show that just a pre-defined set of in-context examples can boost the performance of a visual programming approach (Table 1).
@@
Figure 2: Video Understanding and Reasoning Framework (VURF) pipeline. Top: figure shows the main approach of VURF with the added self-correction module. Bottom: figure shows the self-refinement module.
@@ 3 Methodology
In contrast to conventional task-specific models that exhibit limitations in addressing complex reasoning challenges, the Video Understanding and Reasoning Framework (VURF) seeks to utilize the reasoning power of LLMs to deconstruct complex video-related queries into a series of sub-tasks (video programs). By executing these sub-tasks sequentially, we can culminate to arrive at the final response. Moreover, VURF allows seamless integration of new visual models in a plug-and-play manner and also employs self-critique mechanisms to mitigate LLM’s hallucinations and judgment errors. An overview of our approach is shown in Fig. 2, and we explain the approach in detail below.
@@ 3.1 Video Reasoning LLMs
Role of Large Language Models. Large language models (LLMs), exemplified by GPT-3 Brown et al. (2020) and GPT-4 Achiam et al. (2023), have demonstrated an impressive in-context learning capability (ICL) that does not necessitate fine-tuning. ICL aims to extend the understanding of LLMs to novel scenarios using a restricted set of input and output demonstrations within the relevant context. In this work, we leverage GPT-3.5 to generate visual programs that solve video reasoning tasks involving natural language instructions. The LLMs efficiently perform video reasoning tasks by avoiding direct video processing. Instead, it formulates logical sequence flows that decompose complex tasks into simpler sub-tasks. This approach enhances efficiency and allows the model to navigate the intricacies of video-related challenges.
Prompting. We prompt GPT-3.5 with in-context examples which consist of pairs of instruction and the associated programs that the LLM is expected to generate. The programs follow a generic structure where each line of the program includes the name of a module, the module’s input argument names and their values, and an output variable name. As output variables in a specific step are used later for another step they follow a general structure which the LLM learns:
@@ 3.2 Self-Refinement
One prominent limitation of a naive visual programming technique like Gupta and Kembhavi (2023) is the proneness of generating inaccurate information influenced by contextual cues, as well as lacking an inherent capacity for autonomous self-correction through task-agnostic knowledge. Two major issues arise with the LLM-generated program. Firstly, due to LLM hallucinations, the program might be using some function that is not supported by our interpreter. Secondly, the program may not break down the instruction into sub-tasks in an optimal fashion. We resolve these limitations by the below steps.
@@
Auto-Refinement of In-Context Examples: The effectiveness of our approach hinges significantly on the quality of in-context examples. Therefore, it is imperative to refine these examples to enhance the module’s performance Lu et al. (2023); Madaan et al. (2023). To accomplish this, we implement a self-refinement procedure. Initially, given an instruction, $I$, and the initial program generated $P$, we input $I$ to the LLM, prompting it to generate an improved program, $P^{\prime}$, without the inclusion of in-context examples. Subsequently, both $P$ and $P^{\prime}$ are input into the LLM, enabling the generation of a refined program that aligns more closely with the LLM’s reasoning, all while excluding the influence of in-context examples. Replicating this process for $n$ instructions yields a set of $n$ new in-context examples, enhancing the module’s ability to perform tasks effectively. Note that as shown in Fig. 2, the auto self-refinement module can be applied to a single user query to iteratively improve the generated program as well. However, this is inefficient and costly and thus we show that even pre-refining the in-context program-instruction pairs can improve the performance of the VURF (Table 1). A concrete illustration of this process on a single program and instruction pair is provided in Fig. 3.
Figure 3: Auto Self-Refinement example. Two programs are generated: one with contextual examples and one without, but with added information for structural integrity. Both are then input into the Language Model (LLM) to generate a new program that aligns with the ideal while avoiding invalid functions.
@@ 4 Tasks
Our video understanding and reasoning framework (VURF) aims to offer a versatile approach adaptable to various visual tasks. By integrating an interpreter component into an existing state-of-the-art (SOTA) vision model, VURF addresses four diverse challenges: Video Question Answering (VQA), Video Anticipation, Pose Estimation, and Multi-Video VQA. VURF functions by employing a large language model (LLM) as a reasoning module to generate a visual program. This program outlines a sequence of steps, each executed independently. The output of one step is fed as input to the next, creating a cohesive workflow.